# Perspective on Moreau–Yosida Regularization in Density-Functional Theory

### v1

@@ VI.1 Hartree auto-regularization
Through a special choice for the potential space previously mentioned in Sections III.2 and III.3, we can achieve automatic regularization of the energy functional in the Hartree approximation of DFT. This choice is the homogeneous Sobolev space [75; 76],
$X^{*}=\{u\in L^{2}(\mathbb{R}^{3})\mid\nabla u\in L^{2}(\mathbb{R}^{3}:\mathbb{R}^{3})\},$ (100)
@@
Thus we have found an explicit expression for the duality map, $J(x)=v_{\mathrm{H}}(f)$, when $x\in X$ is a regular distribution represented by $f\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{3})$. Note that we usually use the symbols for distributional densities ‘$x$’ and their representing functions ‘$f$’ for elements of $X$ interchangeably, so we could have also written $J(x)=v_{\mathrm{H}}(x)$. Now take the norm in $X$ with one-half in front,
@@
which retrieves the Hartree mean-field energy from Eq. (42) that is also the energy of the field induced by the charge distribution $f\in X$. Through these identifications, it becomes apparent that this choice of density and potential space is a very natural one for DFT and we believe that it can have a future importance for the formulation of DFT. Indeed, the use of these spaces in DFT was already suggested around 2012, based on an appearance in regularization of optimized-effective-potential methods [77], but never published. Still, several important mathematical questions are left to be studied in this setting, especially the self-adjointness and boundedness below of the Hamiltonian and lower semicontinuity of the ensemble constrained-search functional (Assumption 4). It is expected that the last point only holds on a bounded domain or through including a confining offset potential $v_{\mathrm{o}}$.
@@
which becomes strictly convex through the addition of the Hartree mean-field energy. We are now prepared to show the auto-regularization of the corresponding energy functional from the Legendre–Fenchel transform (the attentive reader might already have noticed that a strictly convex functional $F^{\rm Hartree}(\rho)$ implies a MY-regularized convex conjugate functional). With $v=J(x)$ we get
@@
We can note several things here. In the third line we recognize that the considered energy content of the system is the (non-interacting) kinetic energy $F^{0}(\rho)$ of the system, the potential energy from coupling to the external $v$, but also the energy of the induced field $E_{\rm H}(\rho)$. This sums up to give the self-interaction error of the Hartree approximation. The regularization parameter can be read off directly from the fifth line of Eq. (107) as $\varepsilon=1$. Without including the vacuum permittivity into the norm we can have $\varepsilon=\varepsilon_{0}$, but then the duality map does not directly map densities to potentials any more. The new argument of the regularized functional is $x=-J^{-1}(v)=\varepsilon_{0}\Delta v$ and can of course be negative as a general element in $X$. Note that now the energy functional $E^{\rm Hartree}(v)$ is the result of MY regularization, not the universal density functional as before. Even though we subtract $\frac{1}{2}\|v\|_{X^{*}}^{2}$, which reminds of the way MY regularization was introduced in Section III.2, the energy functional does not necessarily become strictly concave (yet it is concave) since $F^{0}_{\varepsilon}(-x)$ is convex. Since $\rho=\nabla E^{\rm Hartree}(v)$ is now unique, there arises no ambiguity in the ground-state density.
@@ VI.2 Maxwell–Schrödinger DFT
Maxwell–Schrödinger DFT [78] is an extension of standard DFT with electromagnetic fields and internal currents that also considers induced magnetic fields. In this respect it is exactly equivalent to the Hartree method discussed before in Section VI.1, adding a current-current integral instead of the density-density integral of Eq. (105). Since the currents couple to an external vector potential, we include those into the description, just like in current DFT [42]. The vector potentials live in a special homogeneous and divergence-free (Coulomb gauged) Sobolev space that is similar to the previously mentioned homogeneous Sobolev space,
@@
It was possible to rewrite this as a set of component-wise Poisson equations since in the application of the vector Laplacian $\Delta\nabla(\nabla\cdot)-\nabla\times(\nabla\times)$ the divergence term vanishes for $\in Y^{*}$. The current space $Y$ includes distributional vector fields and the norm for regular distributions $\bm{j}\in Y$ features a current-current integral,
@@
In contrast to the original publication [78], we do not include spin degrees-of-freedom, since this makes the notation more economic and does not change the main argument about auto-regularization. We now define a Hamiltonian $\hat{H}^{\lambda=1}_{v,}$ like in Eq. (23) just replacing $\hat{T}$ with $\hat{T}$. We further always choose the fully interacting case with $\lambda=1$ here and will thus suppress this parameter in the notation. The ground-state energy for a potential pair $(v,)\in X^{*}\times Y^{*}$ is then
$E(v,)=\inf_{\Psi}\langle\hat{\Psi},H_{v,}\Psi\rangle,$ (112)
which can equivalently be defined with variation over density matrices. This shows that this functional is again concave in $v$, but has no special convexity property in , a fact related to diamagnetism [78]. We define the corresponding constrained-search functional over ensembles [79; 80; 81],
@@
One immediately sees that the density now additionally couples to $\tfrac{1}{2}^{2}$ which means the resulting energy functional is also not concave in . This can easily be repaired by substituting $u=v+\tfrac{1}{2}^{2}$ which also makes the energy functional jointly concave in $u$ and . Yet note that this construction requires $^{2}\in X^{*}$, the space of potentials, which only holds when the spaces $X$ and $Y$ have a special compatibility property [15].
@@
In this step we introduced a new magnetic coupling constant $\chi$. Now it turns out that the last two functionals, ${\bar{e}}_{\rm M}(v,\bm{j}^{\mathrm{p}})$ and ${\bar{f}}_{\rm M}(\rho,\bm{j}^{\mathrm{p}})$, can be given in the form of Lasry–Lions regularizations [82], a form of double MY regularization, and that by a theorem due to Attouch and Aze [83] they both become Fréchet differentiable if $\chi>1$. This differentiability allows for a unique mapping from paramagnetic currents to vector potentials $\bm{A}$, either for given $v$ or $\rho$, and thus a form of a Hohenberg–Kohn theorem in the context of Maxwell–Schrödinger DFT. The details of this procedure can be found in the original publication by Tellgren [78].
If we remember that the Coulomb potential arises from a QED treatment of the longitudinal component of the electric field in Coulomb gauge [84], then the relation between the Hartree approximation (with only mean-field interaction) and standard DFT is the same as between Maxwell–Schrödinger DFT and a full (non-relativistic) QED treatment in terms of a functional theory. The development of such a theory, called QEDFT, is an ongoing program [85; 86; 87; 24; 88] with important implications for fields like polaritonic chemistry [89], and it might greatly benefit from MY regularization on its theoretical side.
@@ VII Application of MY-based Kohn–Sham inversion to periodic settings
A key advantage of the periodic setting for exploration is that natural choices for the density and potential spaces, $X$ and $X^{*}$, are the periodic Sobolev spaces $H_{\mathrm{per}}^{s}(\Omega)$ with $\Omega\subset\mathbb{R}^{3}$ denoting the periodically repeated unit cell [16, Sec. 2]. These spaces are Hilbert spaces (thus fulfill Assumption 3) and are easily characterized mathematically using Fourier modes $e_{\bm{G}}=e^{\mathrm{i}\bm{G}\cdot\bm{r}}/\sqrt{|\Omega|}$, namely by
$H_{\mathrm{per}}^{s}(\Omega,\mathbb{C}):=\left\{u=\sum_{\bm{G}}\hat{u}_{\bm{G}}e_{\bm{G}}\middle|\left\|u\right\|_{H_{\mathrm{per}}^{s}}<\infty\right\}$ (121)
@@
and where the sum is over the usual reciprocal lattice vectors $\bm{G}$. This characterization is very much in line with the usual discretization schemes based on plane waves, where essentially one just selects a subspace of $H_{\mathrm{per}}^{s}(\Omega)$ by limiting the largest $|\bm{G}|$. Note that the space with $s=+1$ is different to the homogeneous Sobolev space of Eq. (100), since the $1+|\bm{G}|^{2}$ (instead of just $|\bm{G}|^{2}$) in the definition of the norm in Eq. (122) takes care of lower-order derivatives as well.
As a result of this definition, key quantities of the MY program, such as the duality map $J$, can be obtained in computable, closed-form expressions. For example, with the particular choice $X=H_{\mathrm{per}}^{-1}(\Omega)$ for densities and $X^{*}=H_{\mathrm{per}}^{1}(\Omega)$ for potentials one obtains the duality map in the Fourier representation as [19]
@@
Noticing that $1/(\mu+\left|\bm{G}\right|^{2})$ is the Fourier representation of the Yukawa potential $\Phi^{\mu}(\bm{r})=\exp{(-\mu\left|\bm{r}\right|)}/(4\pi\left|\bm{r}\right|)$ then admits the closed-form formulation as a convolution,
@@
Targeting KS inversion in this setting we (as usual) assume that we are given a simultaneously interacting and non-interacting $v$-representable ground-state density $\rho_{\mathrm{gs}}$, meaning that there exists a KS potential $v_{\mathrm{s}}\in-\partial F^{0}(\rho_{\mathrm{gs}})$ and a potential $v\in-\partial F^{1}(\rho_{\mathrm{gs}})$. While $F^{0}(\rho)$ could be directly subjected to the MY program, results seem to generally improve if a maximal amount of known information about the problem is explicitly employed [2; 67]. We therefore choose a ‘guiding functional’
@@
    an approach to compute the proximal points $\rho_{\mathrm{gs,\varepsilon}}$ and
@@
    a suitable sequence of decreasing values for $\varepsilon$ enabling an extrapolation of the $\varepsilon\to 0$ limit.
@@
with respect to the orbitals $\phi_{i}$ and occupations $f_{i}$. In these expressions we do not make explicit any details about spin or Brillouin zone sampling, i.e., the choice of the $\bm{k}$-point grid to model periodic supercells [90, Ch. 2.8]. In that sense one could think of the index $i$ to run over all occupied orbitals of all spins and all reducible $\bm{k}$-points.
@@
We remark that this energy expression Eq. (130) is closely related to the usual KS-DFT energy expression, where essentially the exchange–correlation energy term is just replaced by the penalty term of the regularization. Standard direct minimization techniques from KS-DFT are therefore directly applicable [91; 92; 93]. A first implementation of this approach in DFTK [19; 94] is capable to perform proximal-point computations for on non-trivial 3D solid-state systems, such as silicon, gallium arsenide or potassium chloride, providing thus a practical MY-based KS inversion procedure. Right now the treatment is still restricted to insulating solids, since the algorithm currently makes the assumption that all $f_{i}=1$ in Eq. (129), respectively only minimizes over the $N$ (spin) orbitals in Eq. (130). A generalization to $f_{i}\neq 1$, required for the treatment of metals, is possible taking inspiration from the KS-DFT literature [92; 95; 96; 97].
@@
When minimizing the energy functional of usual KS problems, it is primarily the kinetic-energy term that deteriorates the conditioning due to its large-$\bm{G}$ divergence. Appropriate preconditioners are available to tame this behavior [98; 92]. While for large $\varepsilon$ values these are appropriate for the MY proximal point computation, for small $\varepsilon$ the penalty term $\frac{1}{2\varepsilon}\|\rho_{\Phi}-\rho_{\mathrm{gs}}\|_{H_{\mathrm{per}}^{-1}}^{2}$ can dominate. Refined preconditioning techniques for the MY setting that are $\varepsilon$-dependent and take both terms into account therefore need to be developed.
@@
Figure 3: Convergence of $\rho$ and $v_{\mathrm{xc},\varepsilon}$ as a function of $\varepsilon$ for the bulk silicon test case and a basis truncation of $E_{\text{cut}}=30$, leading to an error $\Delta\rho=2.7\times 10^{-5}$. We refer the reader to the original reference [19] for details of the computational setup.
@@
An additional complication in practical inverse KS calculations is that the input density $\rho_{\mathrm{gs}}$ itself is usually obtained from an inexact procedure, such as an experiment or a computation from a higher-level theory. As a result we do not actually have access to the exact $\rho_{\mathrm{gs}}$, but only a perturbed version $\tilde{\rho}_{\text{gs}}=\rho_{\mathrm{gs}}+\delta\rho$. Computing the proximal point for a given $\varepsilon$ with this perturbed ground state density thus only provides us with an estimate ${\tilde{v}}_{\mathrm{xc},\varepsilon}$ instead of the correct $v_{\mathrm{xc},\varepsilon}$. Previous analysis [19] lets us deduce the bound
@@
for the error between the numerically accessible ${\tilde{v}}_{\mathrm{xc},\varepsilon}$ and the desired true $v_{\mathrm{xc}}$. We note the two error contributions, where the first is controlled by taking $\varepsilon\to 0$. In contrast, the second term due to the inherent error $\delta\rho$ in the available ground-state density data will increase with decreasing $\varepsilon$. Therefore one usually obtains a convergence behavior similar to Figure 3 for the density and potential. Notably the potential $v_{\mathrm{xc},\varepsilon}$ initially converges towards the true $v_{\mathrm{xc}}$, but then diverges, suggesting that there is an optimal $\varepsilon$ value to employ to yield the most accurate estimate of the inverted potential.
@@
The bound in Eq. (131) provides a qualitative explanation of the observed behavior as $\varepsilon\to 0$. However, it is neither sharp nor does it fully match the practical setting of periodic calculations, such that it is not yet sufficient to suggest an optimal $\varepsilon$ value a priori, let alone an entire $\varepsilon$ sequence for potential extrapolation. For example, for efficient plane-wave calculations employing a pseudopotential approximation is required. This implies the addition of a non-local (Kleinman–Bylander) term [99, Ch. 7] to the energy functional Eq. (130), which is outside of the scope of the current mathematical framework [19].
@@
While the current analysis thus already provides first insights into the optimal choice of an $\varepsilon$ sequence (ingredient ii), it still requires considerable refinement until an a priori mathematically guided selection of such a sequence is feasible. An explicit numerical study of the $\varepsilon\to 0$ limit, e.g., based on a logarithmically spaced $\varepsilon$ grid, is thus generally required to empirically identify the optimal $\varepsilon$. However, the progress towards first analytical error bounds such as Eq. (131) as well as the rich literature on a priori and a posterior error analysis for the closely related problem of periodic KS-DFT [100; 101; 102; 103; 104; 105; 106; 107] make it hopeful that more refined and comprehensive analysis is feasible. A tantalizing outlook would be sharper error bounds, thus a better estimation of an optimal $\varepsilon$, or the combination of convergence properties of ${\tilde{v}}_{\mathrm{xc},\varepsilon}$ with extrapolation techniques to accelerate convergence.

### v2

@@ VI.1 Hartree auto-regularization
Through a special choice for the potential space previously mentioned in Sections III.2 and III.3, we can achieve automatic regularization of the energy functional in the Hartree approximation of DFT. This choice is the homogeneous Sobolev space [68; 34; 70],
$X^{*}=\{u\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{3})\mid\nabla u\in L^{2}(\mathbb{R}^{3}:\mathbb{R}^{3})\},$ (100)
@@
Thus we have found an explicit expression for the duality map, $J(x)=v_{\mathrm{H}}(f)$, when $x\in X$ is a regular distribution represented by $f\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{3})$. Note that we usually use the symbols for distributional densities ‘$x$’ and their representing functions ‘$f$’ for elements of $X$ interchangeably, so we could have also written $J(x)=v_{\mathrm{H}}(x)$. Further note that since the potentials are defined modulo a constant it must also hold $\langle u,x\rangle=\langle u+c,x\rangle$ which implies $\int f(\bm{r})\mathrm{d}\bm{r}=0$, i.e., all integrable $f$ representing an $x\in X$ must integrate to zero. This may seem problematic if we think of densities $f$ as normalized to a certain particle number, but it can be compensated by the introduction of a fixed reference density and shifting $X$ to an affine space with the same dual $X^{*}$. Now take the norm in $X^{*}$ with one-half in front,
@@
which retrieves the Hartree mean-field energy from Eq. (42) that is also the energy of the field induced by the charge distribution $f\in X$. Through these identifications, it becomes apparent that this choice of density and potential space is a very natural one for DFT and we believe that it can have a future importance for the formulation of DFT. Indeed, the use of similar spaces in DFT was already suggested around 2012, based on an appearance in regularization of optimized-effective-potential methods [41], but never published. Still, several important mathematical questions are left to be studied in this setting, especially the self-adjointness and boundedness below of the Hamiltonian and lower semicontinuity of the ensemble constrained-search functional (Assumption 4). Like the summary of Section II.4 states, to make the the last point even possible to hold together with Assumption 1, one could switch to a bounded domain or include a confining offset potential $v_{\mathrm{off}}$.
@@
which becomes strongly (and thus strictly) convex through the addition of the Hartree mean-field energy. We are now prepared to show the auto-regularization of the corresponding energy functional from the Legendre–Fenchel transform (the attentive reader might already have noticed that a strictly convex functional $F^{\rm Hartree}(\rho)$ implies a MY-regularized convex conjugate functional). With $v=J(x)$ we get
@@
We can note several things here. In the third line we recognize that the considered energy content of the system is the (non-interacting) kinetic energy $F^{0}(\rho)$ of the system, the potential energy from coupling to the external $v$, but also the energy of the induced field $E_{\rm H}(\rho)$. This sums up to give the self-interaction error of the Hartree approximation. The regularization parameter can be read off directly from the fifth line of Eq. (107) as $\varepsilon=1$. Without including the vacuum permittivity into the norm we can have $\varepsilon=\varepsilon_{0}$, but then the duality map does not directly map densities to potentials any more. The new argument of the regularized functional is $x=-J^{-1}(v)=\varepsilon_{0}\Delta v$ and can of course be negative as a general element in $X$. Note that now the energy functional $E^{\rm Hartree}(v)$ is the result of MY regularization, not the universal density functional as before. Even though we subtract $\frac{1}{2}\|v\|_{X^{*}}^{2}$, which reminds of the way MY regularization was introduced in Section III.2, the energy functional does not necessarily become strictly concave (yet it must be concave by construction) since $F^{0}_{\varepsilon}(-x)$ is convex. Since $\rho=\nabla E^{\rm Hartree}(v)$ is now unique, there arises no ambiguity in the ground-state density.
@@ VI.2 Maxwell–Schrödinger DFT
Maxwell–Schrödinger DFT [98] is an extension of standard DFT with electromagnetic fields and internal currents that also considers induced magnetic fields. In this respect it is similar to the Hartree method discussed before in Section VI.1, adding a current-current integral instead of the density-density integral of Eq. (105). Since the currents couple to an external vector potential, we include those into the description, just like in current DFT [78]. The vector potentials live in a special homogeneous and divergence-free (Coulomb gauged) Sobolev space that is similar to the previously mentioned homogeneous Sobolev space,
@@
It was possible to rewrite this as a set of component-wise Poisson equations since in the application of the vector Laplacian $\Delta\nabla(\nabla\cdot)-\nabla\times(\nabla\times)$ the divergence term vanishes for $\in Y^{*}$ due to the chosen Coulomb gauge. The current space $Y$ includes distributional vector fields and the norm for regular distributions $\bm{j}\in Y$ features a current-current integral,
@@
In contrast to the original publication [98], we do not include spin degrees-of-freedom, since this keeps the notation more concise but does not change the main argument about auto-regularization. We now define a Hamiltonian $\hat{H}^{\lambda=1}_{v,}$ like in Eq. (23) just replacing $\hat{T}$ with $\hat{T}$. We further always choose the fully interacting case with $\lambda=1$ here and will thus suppress this parameter in the notation. The ground-state energy for a potential pair $(v,)\in X^{*}\times Y^{*}$ is then
$E(v,)=\inf_{\Psi}\langle\Psi,\hat{H}_{v,}\Psi\rangle,$ (112)
which can equivalently be defined with variation over density matrices. This shows that this functional is again concave in $v$, but has no special convexity property in , a fact related to diamagnetism [98]. We define the corresponding constrained-search functional over ensembles [105; 106; 57],
@@
One immediately sees that the density now additionally couples to $\tfrac{1}{2}^{2}$, which means that the resulting energy functional is also not concave in . This can easily be repaired by substituting $u=v+\tfrac{1}{2}^{2}$, which makes the energy functional jointly concave in $u$ and . But note that this construction requires $^{2}\in X^{*}$, the space of potentials, which only holds when the spaces $X$ and $Y$ have a special compatibility property [56].
@@
In this step we introduced a new magnetic coupling constant $\chi$. Now it turns out that the last two functionals, ${\bar{e}}_{\rm M}(v,\bm{j}^{\mathrm{p}})$ and ${\bar{f}}_{\rm M}(\rho,\bm{j}^{\mathrm{p}})$ can be given in the form of Lasry–Lions regularizations [60], a form of double MY regularization, that has the formula
$f_{\varepsilon,\delta}(x)=\sup_{z\in X}\inf_{y\in X}\left\{f(y)+\frac{1}{2\varepsilon}\|x-y\|_{X}^{2}-\frac{1}{2\delta}\|x-z\|_{X}^{2}\right\}.$ (121)
By a theorem due to Attouch and Aze [2] they both become Fréchet differentiable if $\chi>1$. This differentiability allows for a unique mapping from paramagnetic currents to vector potentials $\bm{A}$, either for given $v$ or $\rho$, and thus gives a form of Hohenberg–Kohn theorem in the context of Maxwell–Schrödinger DFT. The details of this procedure can be found in the original publication by Tellgren [98].
If we remember that the Coulomb potential arises from a QED treatment of the longitudinal component of the electric field in Coulomb gauge [40], then the relation between the Hartree approximation (with only mean-field interaction) and standard DFT is the same as between Maxwell–Schrödinger DFT and a full (non-relativistic) QED treatment in terms of a functional theory. The development of such a theory, called QEDFT, is an ongoing program [87; 86; 65; 6; 5] with important implications for fields like polaritonic chemistry [88], and it might greatly benefit from MY regularization on its theoretical side.
@@ VII Application of MY-based Kohn–Sham inversion to periodic settings
A key advantage of the periodic setting for exploration is that natural choices for the density and potential spaces, $X$ and $X^{*}$, are the periodic Sobolev spaces $H_{\mathrm{per}}^{s}(\Omega)$ with $\Omega\subset\mathbb{R}^{3}$ denoting the periodically repeated unit cell [18, Sec. 2]. These spaces are Hilbert spaces (thus fulfill Assumption 3) and are easily characterized mathematically using Fourier modes $e_{\bm{G}}=e^{\mathrm{i}\bm{G}\cdot\bm{r}}/\sqrt{|\Omega|}$, namely by
$H_{\mathrm{per}}^{s}(\Omega)=\left\{u=\sum_{\bm{G}}\hat{u}_{\bm{G}}e_{\bm{G}}\middle|\hat{u}_{\bm{G}}^{*}=\hat{u}_{-\bm{G}},\left\|u\right\|_{H_{\mathrm{per}}^{s}}<\infty\right\}$ (122)
@@
where the sum is over the usual reciprocal lattice vectors $\bm{G}$. This characterization is very much in line with the usual discretization schemes based on plane waves, where essentially one just selects a subspace of $H_{\mathrm{per}}^{s}(\Omega)$ by limiting the largest $|\bm{G}|$. Note that the space with $s=+1$ is different to the homogeneous Sobolev space of Eq. (100), apart from consisting of periodic functions, since the $1+|\bm{G}|^{2}$ (instead of just $|\bm{G}|^{2}$) in the definition of the norm in Eq. (123) means the uniform $\bm{G}=0$ component is included as well.
As a result of the above definition of the norm in terms of a Fourier expansion, key quantities of the MY program, such as the duality map $J$, can be obtained in computable, closed-form expressions. For example, with the particular choice $X=H_{\mathrm{per}}^{-1}(\Omega)$ for densities and $X^{*}=H_{\mathrm{per}}^{1}(\Omega)$ for potentials one obtains the duality map in the Fourier representation as [43]
@@
Noticing that $1/(\mu^{2}+\left|\bm{G}\right|^{2})$ is the Fourier representation of the Yukawa potential $\Phi^{\mu}(\bm{r})=\exp{(-\mu\left|\bm{r}\right|)}/(4\pi\left|\bm{r}\right|)$ then admits the closed-form formulation as a convolution,
@@
Targeting KS inversion in this setting we (as usual) assume that we are given a simultaneously interacting and non-interacting $v$-representable ground-state density $\rho_{\mathrm{gs}}$, meaning that there exists a KS potential $v_{\mathrm{s}}\in-\partial F^{0}(\rho_{\mathrm{gs}})$ and a potential $v\in-\partial F^{1}(\rho_{\mathrm{gs}})$. While $F^{0}(\rho)$ could be directly subjected to the MY program, results seem to generally improve if a maximal amount of known information about the problem is explicitly employed [112; 91]. We therefore choose a ‘guiding functional’
@@
    an approach to compute the proximal points $\rho_{\mathrm{gs,\varepsilon}}$, which we will discuss next, and,
@@
    a suitable sequence of decreasing values for $\varepsilon$ enabling an extrapolation of the $\varepsilon\to 0$ limit. Despite progress in the development of analytical error bounds for MY-based KS inversion (MYiKS), the choice of such an $\varepsilon$-sequence remains empirical to date. An explicit convergence study is thus still required as we will outline below.
@@
with respect to the orbitals $\phi_{i}$ and occupations $f_{i}$. In these expressions we do not make explicit any details about spin or Brillouin zone sampling, i.e., the choice of the $\bm{k}$-point grid to model periodic supercells [64, Ch. 2.8]. In that sense one could think of the index $i$ to run over all occupied orbitals of all spins and all reducible $\bm{k}$-points.
@@
We remark that this energy expression Eq. (131) is closely related to the usual KS-DFT energy expression, where essentially the exchange–correlation energy term is just replaced by the penalty term of the regularization. Standard direct minimization techniques from KS-DFT are therefore immediately applicable [72; 52; 28]. A first implementation of this approach in DFTK [43; 44] is capable of performing proximal-point computations for non-trivial 3D solid-state systems, such as silicon, gallium arsenide or potassium chloride, thus providing a practical MYiKS procedure. Right now the treatment is still restricted to insulating solids, since the algorithm currently makes the assumption that all $f_{i}=1$ in Eq. (130) and respectively only minimizes over the $N$ (spin) orbitals in Eq. (131). A generalization to $f_{i}\neq 1$, required for the treatment of metals, is possible taking inspiration from the KS-DFT literature [52; 67; 33; 39].
While the current practical MYiKS scheme is still far from optimized in terms of employed numerical algorithms and implementation, we want to provide a high-level idea of their computational cost. Similar to standard KS-DFT, the main computational bottlenecks of MYiKS are the matrix-vector products involving the gradient of $\mathcal{E}$, which is an operator similar to the usual KS Hamiltonian. For plane-wave basis sets a gradient-vector product scales as $O(N_{b}\log N_{b})$, where $N_{b}$ is the number of plane waves. One iteration step in the minimization of Eq. (131) typically requires one gradient-vector product. With current algorithms a few tens (for large $\varepsilon$) to a few hundred (for small $\varepsilon$) of minimization steps are required to compute one proximal density. The advantage of the MYiKS approach is therefore that no explicit diagonalization of a KS-like Hamiltonian is required—in contrast to for example the ZMP [112] or the Wu–Yang [109] methods. However, the disadvantage of current MYiKS techniques are (i) the sometimes large number of minimization steps per proximal point computation and (ii) the need to estimate the limit $\varepsilon\to 0$, i.e., the need to perform multiple proximal point computations for different $\varepsilon$ values. We now describe some ideas to overcome these challenges.
@@
Crucial to reduce the number of minimization steps is an appropriate preconditioning of the minimization of $\mathcal{E}$. For the usual KS problems it is primarily the kinetic-energy term with its $|\bm{G}|^{2}$-dependence that leads to ill-conditioned problems in large plane-wave basis sets. Appropriate preconditioners have been developed to effectively tame this source of ill-conditioning [99; 52] for KS DFT problems. For the MY proximal-point computation we currently rely on the same kinetic-energy preconditioners. For large $\varepsilon$ values this is indeed sufficient. However, for small $\varepsilon$ the penalty term $\frac{1}{2\varepsilon}\|\rho_{\Phi}-\rho_{\mathrm{gs}}\|_{H_{\mathrm{per}}^{-1}}^{2}$ can dominate, such that kinetic-energy preconditioning alone is no longer appropriate. This explains why sometimes our minimization procedure requires hundreds of steps. Refined preconditioning techniques for the MY setting therefore need to be developed. These should be $\varepsilon$-dependent and take both the kinetic-energy and the penalty term into account.
@@
Figure 4: Convergence of $\tilde{\rho}_{\text{gs},\varepsilon}$ and ${\tilde{v}}_{\mathrm{xc},\varepsilon}$ as a function of $\varepsilon$ for the bulk silicon test case and a basis truncation of $E_{\text{cut}}=30$, leading to an error $\delta\rho=2.7\times 10^{-5}$. We refer the reader to the original reference [43] for details of the computational setup and numerical parameters.
@@
An additional complication in practical inverse KS calculations is that the input density $\rho_{\mathrm{gs}}$ itself is usually obtained from an inexact procedure, such as an experiment or a computation from a higher-level theory. As a result, we typically do not have access to the exact $\rho_{\mathrm{gs}}$, but only to a perturbed version $\tilde{\rho}_{\text{gs}}=\rho_{\mathrm{gs}}+\delta\rho$. Computing the proximal point for a given $\varepsilon$ with this perturbed ground-state density thus only provides us with an estimate ${\tilde{v}}_{\mathrm{xc},\varepsilon}$ instead of the correct $v_{\mathrm{xc},\varepsilon}$. Previous analysis [43] lets us deduce the bound
@@
for the error between the numerically computed ${\tilde{v}}_{\mathrm{xc},\varepsilon}$ and the desired true $v_{\mathrm{xc}}$. We notice two error contributions. The last term is due to the inherent error $\delta\rho$ in the available ground-state density data and increases with decreasing $\varepsilon$. We observe (see Figure 4) that the potential ${\tilde{v}}_{\mathrm{xc},\varepsilon}$ initially converges towards the true $v_{\mathrm{xc}}$, but then the distance increases again: there is an optimal $\varepsilon$ value which yields the most accurate estimate of the inverted potential. We remark that on top of the effects of a noisy input density outlined in Eq. (132) and which we studied in Ref. 43, there is an additional contribution to the total error that we have not made explicit here. This is the error due to the proximal density being only converged up to a certain finite tolerance in practical iterative schemes for minimizing $\mathcal{E}$. Including this term consistently in the analysis is work in progress.
@@
The bound in Eq. (132) provides a qualitative explanation of the observed behavior as $\varepsilon\to 0$. However, it is neither sharp nor does it fully match the practical setting of periodic calculations, such that it is not yet sufficient to suggest an optimal $\varepsilon$ value a priori, let alone an entire $\varepsilon$ sequence for potential extrapolation. For example, while Eq. (132) lets us observe that $\varepsilon$ should not be taken smaller than the error $\delta\rho$, this error is usually not known in practice. Furthermore, the use of a pseudopotential approximation is required for efficient plane-wave calculations. This implies the addition of a non-local (Kleinman–Bylander) term [20, Ch. 7] to the energy functional Eq. (131), which is outside of the scope of the current mathematical framework [43].
@@
While the current analysis thus already provides first insights into the optimal choice of an $\varepsilon$ sequence (ingredient ii), it still requires considerable refinement until an a priori mathematically guided selection of such a sequence is feasible. To date, an explicit numerical study of the $\varepsilon\to 0$ limit, e.g., based on a logarithmically spaced $\varepsilon$ grid, is thus generally required to empirically identify the optimal $\varepsilon$. However, the progress towards first analytical error bounds such as Eq. (132) as well as the rich literature on a priori and a posterior error analysis for the closely related problem of periodic KS-DFT [17; 18; 16; 45; 19; 29; 13; 21] make it hopeful that a more refined and comprehensive analysis is feasible. A tantalizing outlook would be sharper error bounds and thus a better estimation of an optimal $\varepsilon$, or the combination of convergence properties of ${\tilde{v}}_{\mathrm{xc},\varepsilon}$ with extrapolation techniques to accelerate convergence.